Answer the following, final answer must be reduced to lowest term.


  1. The probability of getting even number if a fair die is rolled.
  2. The probability of getting hearts in a deck of a card.
  3. The probability of getting aces in a deck of card
  4. The probability of getting a black cards in a deck of card
  5. The probability of getting a 2 head in tossing 3 coins
1
Expert's answer
2021-04-29T18:00:28-0400
  1. There are 6 possibilities, 3 are odd and 3 are even. So "\frac{3}{6}=" "\frac{1}{2}"
  2. There are 52 cards in the deck, 13 of them are hearts, so "\frac{13}{52}=" "\frac{1}{4}"
  3. There are 52 cards in the deck, 4 of them are aces, so "\frac{4}{52}=" "\frac{1}{13}"
  4. There are 52 cards in the deck, 26 of them are black, so "\frac{26}{52} =" "\frac{1}{2}"
  5. The sample space has size 23 = 8, and there are 3 possible ways to get 2 heads out of 3 coins (1 and 2, 2 and 3, 1 and 3), so total probability is "\frac{3}{8}".